Solve by factoring: 3x^2+16x=12
Answer by stanbon(75887)

Solve by factoring: 3x^2+16x=12
3x^2 + 16x - 12 = 0
---
3x^2 + 18x -2x -12 = 0
---
Factor:
3x(x+6) - 2(x+6) = 0
(x+6)(3x-2) = 0
x = -6 or x = 2/3
=========================
Cheers,
Stan H.